OpenStudy (anonymous):

The area of a circular trampoline is 112.07 square feet. What is the radius of the trampoline? Round to the nearest hundredth. Answer 3.37 feet 10.59 feet 5.97 feet 35.67 feet

OpenStudy (anonymous):

The area of a circle equation is \[A = \pi*r^2\]

OpenStudy (anonymous):

so if the area is 112.07, that means 112.07 = pi*r^2

OpenStudy (anonymous):

112.07 divided by pi equals 35.673 so 35.673 = r^2

OpenStudy (anonymous):

Can you answer it from there?

OpenStudy (anonymous):

square root it

OpenStudy (anonymous):

yup, so sqrt(35.673) is gross looking but a calculator gets you to 5.97269

OpenStudy (anonymous):

from here you round to the nearest hundredth and get to your answer. done and done